# Variational Autoencoders for Efficient Simulation-Based Inference

### v1

@@ 4.3 The SIR Model
Method Two Moons Bernoulli GLM SIR Budget 10,000 20,000 30,000 10,000 20,000 30,000 10,000 20,000 30,000 GATSBI 0.1213 (± 0.0650) 0.0582 (± 0.0122) 0.1432 (± 0.1161) 3.6381 (± 0.5428) 4.8299 (± 0.6825) 3.9965 (± 1.8510) 2.9470 (± 0.4891) 2.7502 (± 0.7268) 2.1466 (± 0.5625) NPE 0.0254 (± 0.0025) 0.0241 (± 0.0099) 0.0125 (± 0.0045) 1.0371 (± 0.1830) 0.9379 (± 0.1001) 0.8675 (± 0.1945) 0.6611 (± 0.2236) 1.0312 (± 0.1255) 1.1903 (± 0.1007) JANA 0.0060 (± 0.0013) 0.0068 (± 0.0017) 0.0047 (± 0.0009) 3.3232 (± 0.0932) 3.1488 (± 0.0280) 2.9010 (± 0.0878) 0.8495 (± 0.1254) 1.1077 (± 0.0336) 1.1421 (± 0.0392) APT 0.0073 (± 0.0030) 0.0069 (± 0.0030) 0.0037 (± 0.0012) 0.7383 (± 0.0853) 0.4944 (± 0.0663) 0.5808 (± 0.0539) 1.1003 (± 0.0712) 1.1093 (± 0.0311) 1.1291 (± 0.0481) CP-VAE 0.0382 (± 0.0110) 0.0369 (± 0.0088) 0.0276 (± 0.0050) 0.1966 (± 0.1375) 0.1557 (± 0.0663) 0.0942 (± 0.0547) 0.8792 (± 0.3327) 1.0158 (± 0.0566) 0.9990 (± 0.1923) UP-VAE 0.0736 (± 0.0097) 0.0795 (± 0.0151) 0.0544 (± 0.0150) 1.5420 (± 0.4930) 1.4748 (± 0.2623) 1.1160 (± 0.7496) 0.5168 (± 0.1193) 0.6884 (± 0.2044) 0.6487 (± 0.1087)
@@ 5 Discussion
Table 3: Training time for different methods across simulation budgets (mean $\pm$ std. dev. in minutes).
Method Two Moons Bernoulli GLM SIR Budget 10,000 20,000 30,000 10,000 20,000 30,000 10,000 20,000 30,000 GATSBI 117.67 (± 2.43) 164.90 (± 57.10) 249.50 (± 105.05) 118.62 (± 0.18) 179.69 (± 49.96) 253.48 (± 113.38) 227.16 (± 1.44) 241.24 (± 7.58) 250.47 (± 2.22) NPE 7.04 (± 0.21) 14.21 (± 0.04) 20.95 (± 0.37) 12.17 (± 0.17) 24.96 (± 0.64) 37.43 (± 0.22) 20.10 (± 5.45) 29.62 (± 3.74) 44.92 (± 9.24) JANA 5.66 (± 0.10) 9.48 (± 0.12) 14.34 (± 0.11) 7.14 (± 0.05) 12.23 (± 0.07) 19.00 (± 0.19) 22.98 (± 0.32) 39.17 (± 2.48) 61.89 (± 5.41) APT 90.31 (± 8.99) 186.17 (± 13.79) 268.28 (± 32.88) 101.39 (± 7.89) 197.30 (± 9.32) 267.38 (± 21.93) 40.15 (± 3.90) 73.74 (± 6.25) 112.51 (± 2.51) CP-VAE 2.23 (± 0.81) 3.75 (± 0.80) 4.99 (± 0.83) 1.76 (± 0.27) 3.64 (± 0.14) 4.93 (± 0.47) 2.77 (± 0.40) 6.06 (± 1.09) 8.68 (± 0.83) UP-VAE 1.98 (± 0.33) 3.27 (± 0.58) 5.35 (± 1.82) 2.10 (± 0.50) 4.43 (± 0.61) 6.57 (± 1.04) 4.44 (± 0.88) 6.71 (± 1.97) 9.60 (± 1.66)
@@
Posterior estimation accuracy: JANA and APT hold an advantage, as JANA is able to capture complex dependencies between parameters, owing to the autoregressive factorization of the joint posterior as a series of conditional distributions; while APT leverages sequential posterior refinement to enhance accuracy.
Computational efficiency: Adversarial training presents well known challenges, which reflect in relatively high training times for GATSBI. The sequential nature of APT adds computational expense over NPE and CP/UP-VAE. Autoregressive models can be computationally expensive to train and evaluate as estimating each parameter depends on previously estimated parameters – however, JANA is able to closely match NPE’s training times for chosen hyperparameters. CP-VAE and UP-VAE are inherently simple, and do not require external summary statistics, aiding scalability and efficient training (Table 3).
Robustness/stability: GATSBI suffers from the instabilities associated with the adversarial setting (e.g., high MMD std. dev. in Table 2). JANA and APT proved to be very robust, with low std. dev. across test problems, followed by CP/UP-VAE and NPE.
Flexibility/approximating complex posteriors: APT holds an advantage over NPE due to its iterative refinement. JANA is particularly suited for posteriors with dependencies between parameters. However, the flexibility of normalizing flow based approaches (NPE, APT, JANA) is hindered by the invertibility requirement. GATSBI is suitable for complex, high-dimensional problems where other approaches may struggle and learning the posterior implicitly becomes important. Both CP-VAE and UP-VAE are highly flexible and adaptable with respect to the architecture.
@@ 6 Conclusion
This paper introduces two latent variable models, CP-VAE and UP-VAE, for amortized inference in likelihood-free simulation-based inference, with a view on efficiency, scalability, interpretability and robustness. The proposed models are validated against state-of-the-art normalizing flow and adversarial training-based approaches on benchmark test problems. Both CP-VAE and UP-VAE deliver comparable performance to the state-of-the-art while being highly computation efficient and scalable. Future enhancements could involve integrating more sophisticated architectures (e.g., Convolutional Neural Networks) to better capture spatial and temporal dependencies within the data. Exploring diverse prior distributions, such as truncated Gaussian priors, may further improve the models’ ability to approximate complex posteriors.
Acknowledgements
@@
The computations/data handling were enabled by the Berzelius resource provided by the Knut and Alice Wallenberg Foundation at the National Supercomputer Centre and by the National Academic Infrastructure for Supercomputing in Sweden (NAISS) at Chalmers e-Commons at Chalmers, and Uppsala Multidisciplinary Center for Advanced Computational Science (UPPMAX) at Uppsala University, partially funded by the Swedish Research Council through grant agreement nos. 2022-06725 and 2018-05973. AH and PS acknowledge support from the Swedish Research Council through grant agreement nos. 2023-05167 and 2023-05593 respectively.

### v2

@@ 3.2 Posterior Visualization on Multimodal Benchmarks
Method SLCP Distractors Bernoulli GLM Raw Lotka Volterra Budget 10,000 20,000 30,000 10,000 20,000 30,000 10,000 20,000 30,000 GATSBI 5.1993 $\pm$1.2320 7.2478 $\pm$1.2260 3.0631 $\pm$0.3244 3.4369 $\pm$0.8517 4.3479 $\pm$0.5433 3.5950 $\pm$0.8613 7.2684 $\pm$0.5733 5.5183 $\pm$1.7126 4.8858 $\pm$1.2146 NPE 1.2167 $\pm$0.2397 0.7834 $\pm$0.3453 0.6428 $\pm$0.1779 1.1227 $\pm$0.0987 1.1389 $\pm$0.1878 1.1448 $\pm$0.1287 2.6357 $\pm$0.3468 3.8635 $\pm$0.2770 4.3310 $\pm$0.0709 JANA 3.4834 $\pm$1.0817 3.1234 $\pm$0.1769 3.0176 $\pm$0.8765 7.6642 $\pm$1.5789 10.4712 $\pm$0.9031 11.3721 $\pm$0.2927 4.8826 $\pm$0.8294 3.7653 $\pm$0.4288 3.8278 $\pm$0.5220 APT 1.8380 $\pm$0.4448 0.8278 $\pm$0.2373 0.4422 $\pm$0.1428 0.7233 $\pm$0.2131 0.6968 $\pm$0.0826 0.5437 $\pm$0.1333 3.1921 $\pm$0.8198 3.2722 $\pm$0.8169 4.4739 $\pm$0.9642 Simformer 0.0540 $\pm$0.0173 0.0541 $\pm$0.0161 0.0552 $\pm$0.0195 0.2056 $\pm$0.0486 0.0926 $\pm$0.0493 0.0731 $\pm$0.0238 0.9696 $\pm$0.4434 0.8342 $\pm$0.4742 0.9634 $\pm$0.2588 CP-VAE 3.8511 $\pm$1.4076 1.8129 $\pm$0.6628 1.2052 $\pm$0.6557 2.1708 $\pm$0.0391 1.6792 $\pm$0.3563 1.6827 $\pm$0.0000 3.1960 $\pm$0.3824 3.2443 $\pm$0.2162 3.4280 $\pm$0.2793 UP-VAE 1.9762 $\pm$0.3219 1.5244 $\pm$0.1978 1.4353 $\pm$0.1306 1.2034 $\pm$0.6426 1.6952 $\pm$1.0074 1.1553 $\pm$0.8829 4.4367 $\pm$0.7106 6.1012 $\pm$0.5511 6.0285 $\pm$1.1408
(a) Gaussian Mixture, Gaussian Linear, Gaussian Linear Uniform, and SLCP
Method Gaussian Mixture Gaussian Linear Gaussian Linear Uniform SLCP Budget 10,000 20,000 30,000 10,000 20,000 30,000 10,000 20,000 30,000 10,000 20,000 30,000 GATSBI 0.2242 $\pm$0.0296 0.1149 $\pm$0.0549 0.1733 $\pm$0.1082 0.4999 $\pm$0.1595 0.6476 $\pm$0.3050 0.7182 $\pm$0.2839 0.9044 $\pm$0.2353 0.7712 $\pm$0.1483 1.1197 $\pm$0.4795 0.7197 $\pm$0.5598 0.3375 $\pm$0.1473 0.7948 $\pm$0.5992 NPE 0.0604 $\pm$0.0260 0.0418 $\pm$0.0211 0.0554 $\pm$0.0347 0.0105 $\pm$0.0036 0.0074 $\pm$0.0018 0.0082 $\pm$0.0011 0.0258 $\pm$0.0098 0.0117 $\pm$0.0041 0.0121 $\pm$0.0051 0.4021 $\pm$0.0551 0.1632 $\pm$0.0684 0.2333 $\pm$0.0935 JANA 0.0579 $\pm$0.0079 0.0415 $\pm$0.0037 0.0383 $\pm$0.0042 0.5375 $\pm$0.0367 0.5245 $\pm$0.0072 0.5169 $\pm$0.0028 0.5798 $\pm$0.0357 0.5875 $\pm$0.0279 0.6181 $\pm$0.0128 0.3673 $\pm$0.2011 0.2045 $\pm$0.0917 0.1209 $\pm$0.0456 APT 0.0137 $\pm$0.0031 0.0108 $\pm$0.0088 0.0110 $\pm$0.0034 0.0075 $\pm$0.0013 0.0070 $\pm$0.0018 0.0080 $\pm$0.0008 0.0166 $\pm$0.0087 0.0138 $\pm$0.0053 0.0113 $\pm$0.0039 0.1405 $\pm$0.0347 0.0487 $\pm$0.0296 0.0287 $\pm$0.0166 Simformer 0.0009 $\pm$0.0007 0.0009 $\pm$0.0006 0.0007 $\pm$0.0010 0.0092 $\pm$0.0116 0.0037 $\pm$0.0035 0.0037 $\pm$0.0028 0.0014 $\pm$0.0006 0.0015 $\pm$0.0011 0.0008 $\pm$0.0003 0.0323 $\pm$0.0203 0.0025 $\pm$0.0034 0.0021 $\pm$0.0030 CP-VAE 0.1846 $\pm$0.0602 0.2623 $\pm$0.1369 0.2209 $\pm$0.0587 0.0382 $\pm$0.0131 0.0220 $\pm$0.0058 0.0216 $\pm$0.0081 0.0829 $\pm$0.0295 0.0448 $\pm$0.0070 0.0363 $\pm$0.0130 0.4309 $\pm$0.1369 0.2488 $\pm$0.0557 0.1658 $\pm$0.0881 UP-VAE 0.3081 $\pm$0.1534 0.3927 $\pm$0.1208 0.3684 $\pm$0.0876 0.0611 $\pm$0.0207 0.0121 $\pm$0.0036 0.0446 $\pm$0.0025 0.0912 $\pm$0.0143 0.1315 $\pm$0.0293 0.1246 $\pm$0.0187 0.4740 $\pm$0.1183 0.4042 $\pm$0.1181 0.3946 $\pm$0.1852
Table 4: Training time (in minutes) for different methods across simulation budgets (mean $\pm$ std. dev.).
(a) Two Moons, Bernoulli GLM, SIR.
Method Two Moons Bernoulli GLM SIR Budget 10,000 20,000 30,000 10,000 20,000 30,000 10,000 20,000 30,000 GATSBI 117.67 $\pm$2.43 164.90 $\pm$57.10 249.50 $\pm$105.05 118.62 $\pm$0.18 179.69 $\pm$49.96 253.48 $\pm$113.38 227.16 $\pm$1.44 241.24 $\pm$7.58 250.47 $\pm$2.22 NPE 7.04 $\pm$0.21 14.21 $\pm$0.04 20.95 $\pm$0.37 12.17 $\pm$0.17 24.96 $\pm$0.64 37.43 $\pm$0.22 20.10 $\pm$5.45 29.62 $\pm$3.74 44.92 $\pm$9.24 JANA 5.66 $\pm$0.10 9.48 $\pm$0.12 14.34 $\pm$0.11 7.14 $\pm$0.05 12.23 $\pm$0.07 19.00 $\pm$0.19 22.98 $\pm$0.32 39.17 $\pm$2.48 61.89 $\pm$5.41 APT 90.31 $\pm$8.99 186.17 $\pm$13.79 268.28 $\pm$32.88 101.39 $\pm$7.89 197.30 $\pm$9.32 267.38 $\pm$21.93 40.15 $\pm$3.90 73.74 $\pm$6.25 112.51 $\pm$2.51 Simformer 3.96 $\pm$0.04 7.36 $\pm$0.11 10.73 $\pm$0.25 17.87 $\pm$0.18 35.05 $\pm$0.52 52.81 $\pm$0.91 4.24 $\pm$0.29 7.67 $\pm$0.38 10.68 $\pm$0.45 CP-VAE 2.23 $\pm$0.81 3.75 $\pm$0.80 4.99 $\pm$0.83 1.76 $\pm$0.27 3.64 $\pm$0.14 4.93 $\pm$0.47 2.77 $\pm$0.40 6.06 $\pm$1.09 8.68 $\pm$0.83 UP-VAE 1.98 $\pm$0.33 3.27 $\pm$0.58 5.35 $\pm$1.82 2.10 $\pm$0.50 4.43 $\pm$0.61 6.57 $\pm$1.04 4.44 $\pm$0.88 6.71 $\pm$1.97 9.60 $\pm$1.66
(b) SLCP Distractors, Bernoulli GLM Raw, Lotka Volterra.
Method SLCP Distractors Bernoulli GLM Raw Lotka Volterra Budget 10,000 20,000 30,000 10,000 20,000 30,000 10,000 20,000 30,000 GATSBI 74.37 $\pm$2.59 100.59 $\pm$2.39 102.30 $\pm$0.51 114.76 $\pm$1.19 115.22 $\pm$1.43 144.10 $\pm$24.15 78.91 $\pm$6.32 75.73 $\pm$0.70 74.99 $\pm$0.20 NPE 21.92 $\pm$2.41 43.92 $\pm$2.58 64.38 $\pm$7.13 25.39 $\pm$2.09 61.43 $\pm$2.61 85.94 $\pm$6.55 80.67 $\pm$14.61 160.92 $\pm$42.82 188.04 $\pm$29.46 JANA 8.49 $\pm$0.06 13.59 $\pm$0.24 20.56 $\pm$0.12 24.02 $\pm$0.37 44.21 $\pm$0.23 64.50 $\pm$1.01 7.59 $\pm$0.02 13.28 $\pm$0.14 18.55 $\pm$0.08 APT 26.30 $\pm$0.36 49.96 $\pm$2.73 74.31 $\pm$2.54 21.87 $\pm$1.28 40.00 $\pm$1.37 60.94 $\pm$3.00 45.52 $\pm$4.85 96.15 $\pm$13.77 144.33 $\pm$18.33 Simformer 15.98 $\pm$1.22 31.98 $\pm$2.34 47.85 $\pm$1.86 16.96 $\pm$0.51 32.46 $\pm$1.89 48.01 $\pm$1.44 5.0 $\pm$0.34 9.23 $\pm$0.54 13.64 $\pm$0.21 CP-VAE 1.54 $\pm$0.34 3.62 $\pm$0.76 4.63 $\pm$1.52 6.04 $\pm$0.99 14.50 $\pm$0.18 10.36 $\pm$0.20 7.22 $\pm$3.25 14.61 $\pm$1.92 22.87 $\pm$1.84 UP-VAE 1.89 $\pm$0.15 2.85 $\pm$0.39 4.25 $\pm$0.52 3.42 $\pm$0.49 8.64 $\pm$2.80 14.21 $\pm$7.27 2.89 $\pm$0.52 4.93 $\pm$0.10 7.16 $\pm$1.26
(c) Gaussian Mixture, Gaussian Linear, Gaussian Linear Uniform, and SLCP
Method Gaussian Mixture Gaussian Linear Gaussian Linear Uniform SLCP Budget 10,000 20,000 30,000 10,000 20,000 30,000 10,000 20,000 30,000 10,000 20,000 30,000 GATSBI 114.07 $\pm$2.11 111.30 $\pm$0.28 137.46 $\pm$19.16 48.19 $\pm$0.24 48.70 $\pm$0.84 49.12 $\pm$0.49 43.27 $\pm$0.36 66.43 $\pm$5.70 69.10 $\pm$0.15 162.72 $\pm$14.73 196.19 $\pm$119.42 99.39 $\pm$5.63 NPE 19.56 $\pm$2.47 42.88 $\pm$4.54 64.03 $\pm$6.51 12.03 $\pm$0.80 24.69 $\pm$1.63 38.79 $\pm$1.23 16.55 $\pm$0.85 30.90 $\pm$2.77 49.42 $\pm$1.49 42.97 $\pm$6.35 97.70 $\pm$6.66 117.61 $\pm$4.12 JANA 7.19 $\pm$0.03 13.95 $\pm$0.94 17.14 $\pm$0.10 2.21 $\pm$0.03 3.68 $\pm$0.05 5.21 $\pm$0.05 1.82 $\pm$0.02 2.96 $\pm$0.01 4.10 $\pm$0.03 29.34 $\pm$0.60 53.76 $\pm$1.05 76.19 $\pm$5.71 APT 76.27 $\pm$7.02 167.88 $\pm$15.99 240.13 $\pm$22.74 70.47 $\pm$0.88 135.38 $\pm$5.75 203.27 $\pm$5.97 78.61 $\pm$2.11 158.57 $\pm$9.32 219.44 $\pm$11.83 93.40 $\pm$1.26 150.29 $\pm$11.10 216.30 $\pm$21.38 Simformer 3.84 $\pm$0.04 7.15 $\pm$0.09 10.48 $\pm$0.21 18.18 $\pm$0.13 35.46 $\pm$0.17 55.46 $\pm$1.13 17.89 $\pm$0.23 33.47 $\pm$0.40 51.24 $\pm$0.89 4.23 $\pm$0.01 7.67 $\pm$0.13 11.32 $\pm$0.40 CP-VAE 2.37 $\pm$0.81 4.67 $\pm$1.10 5.19 $\pm$0.91 3.11 $\pm$0.58 5.25 $\pm$0.63 6.49 $\pm$1.38 3.36 $\pm$0.59 7.48 $\pm$1.65 8.45 $\pm$2.47 3.85 $\pm$0.62 11.38 $\pm$1.97 19.53 $\pm$2.99 UP-VAE 5.61 $\pm$1.20 6.30 $\pm$1.29 12.96 $\pm$1.41 5.75 $\pm$1.16 9.15 $\pm$1.03 9.16 $\pm$1.74 7.41 $\pm$1.85 14.08 $\pm$1.55 18.64 $\pm$6.57 5.49 $\pm$0.65 12.59 $\pm$1.17 22.69 $\pm$3.23
@@ 4 Discussion
Posterior estimation accuracy: Flow-based (NPE, APT, JANA) and diffusion-based methods (Simformer) typically yield higher accuracy, with Simformer often performing best overall. This is expected, as flow and diffusion frameworks tend to model complex posterior distributions more precisely than variational or GAN-based methods, which can suffer from oversimplified approximations, bound-related errors, and adversarial instability.
Computational efficiency: Adversarial training presents well known challenges, which reflect in relatively high training times for GATSBI. The sequential nature of APT adds computational expense over NPE and CP/UP-VAE. Autoregressive models can be computationally expensive to train and evaluate as estimating each parameter depends on previously estimated parameters – however, JANA is able to closely match NPE’s training times for chosen hyperparameters. Simformer, leveraging transformers, has high training time due to the complexity of transformer training and suffers from slow inference speed due to its iterative denoising process. CP-VAE and UP-VAE have simple yet flexible architectures enabling efficient and fast training (Table 4).
Robustness/stability: GATSBI suffers from the instabilities associated with the adversarial setting (e.g., high MMD std. dev. in Table 2). JANA, APT, Simformer proved to be very robust, with low std. dev. across test problems, followed by CP/UP-VAE and NPE.
C2ST v/s MMD: While the C2ST metric is a popular metric for validating inference models in SBI, it may be difficult to interpret in high dimensions, where one poorly estimated dimension can skew the score and results depend on classifier parameters. In contrast, MMD compares empirical distributions directly without relying on a separate classifier, providing a more robust measure in complex, high-dimensional posterior evaluations.
Flexibility/approximating complex posteriors: Simformer is highly effective overall, leveraging transformers and cross-attention conditioning mechanisms to learn discriminative patterns effectively. APT holds an advantage over NPE due to its iterative refinement. JANA is particularly suited for posteriors with dependencies between parameters. However, the flexibility of normalizing flow based approaches (NPE, APT, JANA) is hindered by the invertibility requirement. GATSBI is suitable for complex, high-dimensional problems where other approaches may struggle and learning the posterior implicitly becomes important. Both CP-VAE and UP-VAE are highly flexible and adaptable with respect to the architecture.
@@ 5 Conclusion
This paper introduces two latent variable models, CP-VAE and UP-VAE, for amortized inference in likelihood-free simulation-based inference, with a view on efficiency, interpretability and robustness. The study compares four generative model classes—VAEs, GANs, flow-based methods, and diffusion models—specifically tailored for simulation-based inference, evaluating them on a standardized benchmark for accuracy and training speed. Both CP-VAE and UP-VAE are able to approximate complex posterior distributions while being highly computational efficient and interpretable.
Future enhancements could involve integrating more sophisticated architectures (e.g., Convolutional Neural Networks) to better capture spatial and temporal dependencies within the data. Additionally, incorporating normalizing flows can provide more flexible posterior approximations, enhancing the model’s ability to represent complex distributions. Exploring a variety of prior distributions, including truncated Gaussian priors, may further improve the model’s capacity to approximate intricate posterior distributions.
6 Acknowledgements
@@
The computations/data handling were enabled by the Berzelius resource provided by the Knut and Alice Wallenberg Foundation at the National Supercomputer Centre and by the National Academic Infrastructure for Supercomputing in Sweden (NAISS) at Chalmers e-Commons at Chalmers, and Uppsala Multidisciplinary Center for Advanced Computational Science (UPPMAX) at Uppsala University. Andreas Hellander and Prashant Singh acknowledge support from the Swedish Research Council through grant agreement nos. 2023-05167 and 2023-05593 respectively.
Supplementary Material
Appendix A Hodgkin-Huxley Model
We evaluate our variational methods on the Hodgkin-Huxley (HH) model [Gloeckler et al., 2024], which simulates neuronal action potentials by modeling membrane voltage under sodium, potassium, and leak currents. The parameter vector $\bm{\theta}=(C_{m},g_{\text{Na}},g_{\text{K}},g_{\text{L}},E_{\text{Na}},E_{\text{K}},E_{\text{L}})$ is sampled from a uniform prior, and the system is run for 200 ms with a brief current injection. Since the true posterior is unknown, we evaluate our parameter inference performance by generating 10,000 samples from each model for a specific parameter configuration (1.437,72.177,24.209,0.154,66.87,81.435,67.606)(1.437,72.177,24.209,0.154,66.87,-81.435,-67.606) and by conducting a posterior predictive check—randomly selecting five posterior draws to simulate new voltage traces. As shown in Figure 5, both models provide reasonable posterior estimates, with the UP-VAE generally exhibiting lower variance, while Figure 6 indicates that the generated voltage trajectories closely match the reference voltage signals. The training time for CP-VAE is $4.7\pm 1.17$ minutes, while UP-VAE trains in $5.03\pm 0.04$ minutes, averaged over 5 independent runs with a budget of 10,000.
(a) CP-VAE
(b) UP-VAE
Figure 5: $H$ Model: Estimated posterior (simulation budget: 10,000). (a) CP-VAE and (b) UP-VAE.
(a) CP-VAE
(b) UP-VAE
Figure 6: $H$ Model: Posterior predictive checks (simulation budget: 10,000). (a) CP-VAE and (b) UP-VAE.
Appendix B Hyperparameters
For JANA and Simformer, we use hyperparameters published in their original works [Radev et al., 2023a, Gloeckler et al., 2024]. The NPE approach adopts normalizing spline flows (NSF) from the sbi toolbox, APT follows the same NSF setup [Tejero-Cantero et al., 2020] but distributes each budget across multiple rounds, dividing total simulations by the number of rounds. For GATSBI, hyperparameters for available tasks were taken directly from [Ramesh et al., 2022], while for other test problems, we adapt them according to each problem’s structural and computational requirements. The configurations for UP-VAE and CP-VAE are described below:
B.1 CP-VAE
In this configuration, common parameters are applied uniformly across all problems, with specific architectures detailed in Table 5. Each network — $q_{\bm{\phi}}(\bm{z}\mid\bm{y},\bm{\theta})$, $p(\bm{z}\mid\bm{y})$, and $p(\bm{\theta}\mid\bm{y},\bm{z})$ — is designed to approximate a multivariate distribution, outputting both a mean and variance vector. Specifically, the last layer in $q_{\bm{\phi}}(\bm{z}\mid\bm{y},\bm{\theta})$ and $p(\bm{z}\mid\bm{y})$ outputs a vector of size $2\times\text{latent\_dim}$, while the last layer in $p(\bm{\theta}\mid\bm{y},\bm{z})$ outputs a vector of size $2\times\text{theta\_dim}$. Leaky ReLU activation functions with a negative slope of 0.1 are used throughout, with a learning rate of $5\times 10^{-4}$, except for the Lotka-Volterra model, which uses $10^{-4}$, and the AdamW optimizer [Loshchilov and Hutter, 2019]. Models are trained with a batch size of 32 (except SLCP, which uses 16, and Lotka-Volterra, which uses 128) for 1,000 epochs, with early stopping triggered if validation performance does not improve for 20 consecutive epochs. Both parameters and the data are normalized using standard scaling, and gradient clipping is set to 3.0. The latent dimension is equal to the parameter dimension, except for the Gaussian Linear and GLM raw models, which use 5, Lotka-Volterra uses 8, SLCP uses 4, SLCP with Distractors uses 8, and Gaussian Linear uses 5. Kaiming normal initialization is used for weights. For each simulation budget, 90% of the data is used for training, with the remaining 10% reserved for validation.
Table 5: Architectural details of the CP-VAE model.
Problem $\bm{q_{\phi}(z\mid y,\theta)}$ $\bm{p(z\mid y)}$ $\bm{p(\theta\mid y,z)}$
Gaussian Linear (64,64,64) (64,64) (64,64,64)
Gaussian Lin. Uniform (64,64,64) (64,64) (64,64,64)
SLCP (200,200,200) (64,64) (200,200)
SLCP + Distractors (200,200,200) (64,64) (200,200,200)
Bernoulli GLM (128,128,128) (128,128) (128,128,128)
Bernoulli GLM Raw (256,256,256,256) (128,128) (256,256,256,256)
Gaussian Mixture (64,64,64) (64,64) (64,64,64)
Two Moons (128,128,128) (128,128) (128,128,128)
SIR (100,100) (100,100) (100,100,100)
Lotka-Volterra (128,128,128,128) (64,64) (128,128,128,128)
HH Model (128,128,128,128) (128,128,128,128) (128,128,128,128)
B.2 UP-VAE
Building on CP-VAE, UP-VAE retains core training hyperparameters while introducing key refinements. The architectures of $q_{\phi}(\bm{z}\mid\bm{y},\bm{\theta})$, $p(\bm{\theta}\mid\bm{y},\bm{z})$, and $p(\bm{y}\mid\bm{z})$ are outlined in Table 6. The loss function weights are adjusted to prioritize the theta decoder and latent space constraint, with data reconstruction set to 0.2, theta loss to 0.4, and KL divergence to 0.4.
Table 6: Architectural details of the UP-VAE Model.
Problem $\bm{q_{\phi}(z\mid y,\theta)}$ $\bm{p(\theta\!\mid\!y,z)}$ $\bm{p(y\!\mid\!z)}$
Gauss. Linear (64,64,64) (64,64,64) (64,64,64)
Gauss. Lin. Uniform (64,64,64) (64,64,64) (64,64,64)
SLCP (200,200,200) (200,200,200) (200,200,200)
SLCP + Distractors (200,100,50,25) (200,100,50,25) (128,128,128)
Bernoulli GLM (128,128,128) (128,128,128) (128,128,128)
Bernoulli GLM Raw (128,128,128,128) (128,128,128,128) (128,128,128,128)
Gaussian Mixture (64,64,64) (64,64,64) (64,64,64)
Two Moons (256,128,64) (256,128,64) (256,128,64)
SIR (100,100,100) (100,100,100) (100,100,100)
Lotka-Volterra (128,128,128,128) (128,128,128) (128,128,128)
HH Model (256,128,64) (256,128,64) (256,128,64)
Appendix C Additional Plots
This section presents additional results for Bernoulli GLM Raw and SIR tasks. As Bernoulli GLM considers raw noisy data, CP-VAE’s prior $p(\mathbf{z}|\mathbf{y})$ is affected, as $\mathbf{y}\in\mathbb{R}^{100}$ contains significant noise, impacting training dynamics. In contrast, UP-VAE’s prior $p(\mathbf{z})$ remains unaffected, leading to better performance (Figure 7). For SIR, both models sufficiently capture the parameters, though UP-VAE exhibits higher variance (Figure 8).
(a) CP-VAE
(b) UP-VAE
Figure 7: Bernoulli GLM Raw: Estimated posterior (simulation budget: 30,000). (a) CP-VAE and (b) UP-VAE.
(a) CP-VAE
(b) UP-VAE
Figure 8: SIR: Estimated posterior (simulation budget: 30,000). (a) CP-VAE and (b) UP-VAE.